I-MPC sider the following model of a closed economy in the short run Total demand is given by the accounting identity I MPC YClosed Economy In the Short Run. I have solved (a) but confused on how to solve the rest.
Answer #1

a) Aggregate demandi Y:C+I+G = 20+ air + Tv -T2) + bo-6,1 *.GI = 20 **1 [Y+ Tro-hy-To-99)+60-brit Get = ao+al (tro-to + (1--r(d) Natweal rate of output : change in output for G 7 = 200 ay La = 1.6 i AY = 1.6*12 = 19.2 Change in output for i da ne nab